# Asking Fast Radio Bursts (FRBs) for More than Reionization History

### v1

@@ 4.3 Impact of Non-uniform Distribution of FRBs and Post-EoR
We introduce the contribution from post-EoR ($0\leq z\leq 6$) and consider three different scenarios – (i) Optimistic, where the contribution from the post-EoR is perfectly modelled and completely removed from the data, (ii) Moderate, where the post-EoR contribution is imperfectly removed from the individual ${\rm DM}(z)$ such that the residual low-$z$ contribution remains within $0.5\sigma_{\rm DM}$ at $z=6$, and (iii) Pessimistic, where the post-EoR contribution is completely present for every FRB measured. Note that these three scenarios are differentiated based on the how accurately we know the ${\rm DM}(\bm{\theta},z<6)$ for every observed LoS $\bm{\theta}$ above $z=6$. The assumption that we either know the redshifts of every FRB or at least can separate EoR FRBs from the post-EoR FRBs confidently is implicit within all the three scenarios and is itself non-trivial. Apart from the fact that the boundary between EoR and post-EoR is still fuzzy, the erroneous/indeterminate FRB redshifts can cause a spill over the boundary. This may bias our estimators and we plan to investigate this in greater detail in future studies.
We simulate a matter density LC within the redshift range $0\leq z\leq 6$ (comoving distance $8542~{}{\rm Mpc}$), and assume that the electrons linearly follow the underlying density contrast. We use the publicly available gadget4444https://wwwmpa.mpa-garching.mpg.de/gadget4/ (Springel et al., 2021) to simulate the dark matter density fields within comoving cubes of volume $[500~{}h^{-1}~{}{\rm Mpc}]^{3}$ at the lower redshifts. We choose the same cosmological parameters as mentioned in the end of §1. We also set the resolution (for Particle-Mesh) of the box to match those in the grizzly simulations. We simulated the particle distribution at $66$ redshifts within the range $0\leq z\leq 7$ roughly equidistant by $200~{}{\rm Myr}$. We generated the density coeval boxes on $600^{3}$ grids and then used them to make the density LC. We use weighted linear interpolation scheme to interpolate the density fields at the desired redshift slices from the redshifts at which the coeval boxes are generated. We cannot simulate a large box spanning the whole range up to $z\leq 6$. Hence, we need to repeat the boxes along the LoS ($z$-axis) while creating the final LC.
The contribution from low-redshifts increases both ${\rm\overline{{DM}}_{EoR}}$ and $\sigma_{\rm DM}(z)$, making the separation between the two different histories more challenging. In Figure 16, $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ for the three different histories starts at different $N_{\rm FRB}^{\rm tot}$. This is because, for the slower history, $z_{\rm mid}$ is larger than that of the other histories, requiring a higher $N_{\rm FRB}^{\rm tot}$ to populate the corresponding $z$-bin with at least two FRBs. $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ converges very quickly for $N_{\rm FRB}^{\rm tot}\approx 10$; however the corresponding dispersion is large and it decreases with increasing $N^{\rm tot}_{\rm FRB}$, as expected. $N_{\rm FRB}^{\rm tot}\geq 20$ is sufficient to distinguish $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ at $3\sigma$ for the optimistic scenario. This lower bound remains the same to distinguish the slower reionization history with the faster one for the moderate scenario. However, it takes nearly $N_{\rm FRB}^{\rm tot}=150$ and $40$ to discern $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ of slower-fiducial and fiducial-faster pairs of histories, respectively. Considering the pessimistic case, it takes $N_{\rm FRB}^{\rm tot}=80,~{}150,~{}600$ to distinguish between faster-slower, slower-fiducial and fiducial-faster pairs of histories, respectively.
Figure 16 also shows the variation of the $\langle\overline{{\rm DM}}_{\rm margin}\rangle$ with $N_{\rm FRB}^{\rm tot}$. Considering the optimistic case, we need roughly $40$ FRBs identified during the EoR to discern slower and faster reionization histories at $3\sigma$ with $\langle\overline{{\rm DM}}_{\rm margin}\rangle$. Whereas the $N_{\rm FRB}^{\rm tot}$ slightly increases to $60$ and $90$ for faster-fiducial and fiducial-slower pairs of histories, respectively. These numbers respectively increase to $90$, $200$ and $300$ for the moderate case as shown in the middle panel. Finally for pessimistic scenario, we need $N_{\rm FRB}^{\rm tot}\approx 220,~{}600,~{}1000$ to discern (at $3\sigma$) slower-faster, faster-fiducial and fiducial-slower pairs of reionization histories, respectively. The numbers we found are realistic, and it would be possible to detect many FRBs during EoR using the upcoming SKA-mid.
5 Discussion
Understanding the epoch of reionization (EoR) is a crucial step in learning about one of the most important eras in the cosmic history, when it transitioned from being devoid of any stars, consisting of cold and neutral gas, to hot, ionized, and teeming with the objects we see today. The first sources are supposed to drive the whole process of reionization, therefore studying the IGM during EoR can be connected with the properties and emergence of the first structures. There are many direct and indirect contemporary probes such as the redshifted 21-cm signal from H i in the IGM during EoR and, Thomson scattering optical depth of the CMB photons, high-$z$ quasar spectra, Ly-$\alpha$ systems at high-$z$. However these probes are limited by their own challenges to date. In this work, we propose to use the dispersion measure (${\rm DM}$) of the fast radio bursts (FRBs) from the high-$z$ to probe the IGM during EoR.
@@ 5 Discussion
The dispersion introduced in the FRB pulse, while it travels through the ionized medium, can be used in probing the electron distribution along the line-of-sight (LoS) during EoR. We demonstrated the use of the sky-averaged ${\rm DM}$ and its derivatives to discern between the different reionization histories. Beyond this, we primarily aim to make use of the sky-averaged and the angular dispersion in the ${\rm DM}$ estimates to extract information on the ionization bubbles during the EoR. Using a toy model (Figure 1) of the binary ionization field (within the range $6\leq z\leq 15$), we see that the ${\rm\overline{{DM}}_{EoR}}(z)$ first increases (starting from the end of EoR) roughly up to the mid-point of the reionization, $z_{\rm mid}$, and then tend to saturate as the ionized regions decreases towards the initial stages of the reionization (Figure 2). The derivative $d{\rm\overline{{DM}}_{EoR}}/dz$ directly traces how fast the reionization progresses (Figure 3). The all-sky variance $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ is larger for the reionization scenario where the bubble sizes are larger and vice-versa (Figure 6). We also compute $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ where $\Xi$ is the angular dispersion (eq. 4) at any redshift. We demonstrated that the contours are elongated or squeezed along $z$ and $\delta\theta$ depending respectively on the varying reionization history (Figure 4) and bubble sizes (Figure 7). The impacts are clearly prominent for the marginalized structure function derivatives (Figures 5 and 8).
@@
We also analyzed a more realistic reionization light-cone (LC) (Figure 10) generated using grizzly 1D radiative transfer code for three different reionization histories ending at the same $z$. The behaviour of ${\rm\overline{{DM}}_{EoR}}$ is the same as in the toy model. We find that the different LoS variance plays a significant role for realistic IGM (Figure 11) and biases the average ${\rm\overline{{DM}}_{EoR}}(z)$ by more than $1\sigma$ as compared to the case when the ${\rm DM}$ is computed using the averaged ionization fraction $\bar{x}_{\rm{H}\,\textsc{ii}}$. The slope of the $d{\rm\overline{{DM}}_{EoR}}/dz$ is directly sensitive to the reionization history (Figure 12). $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ is also dependent on the reionization history as well as indirectly on the IGM morphology (bubble sizes) in an intermingled way. Figure 14 clearly shows that $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ is sensitive to the reionization window as well as the typical bubble sizes. The contours are squeezed along $z$ axis and elongated along $\delta\theta$ axis for faster reionization history which has relatively larger ionized bubbles and vice-versa. Our initial analyses are rather optimistic where we have considered the FRBs to be located uniformly at every grid points in our reionization LC with their redshifts known within an uncertainty of $\Delta z=0.5$. We also assumed that the contribution from the low-redshift ($z<6$) has been perfectly removed from the DM measurements.
@@
We next consider observationally more realistic situation where the FRBs abundances vary with redshift (Figure 15) and they are more clustered at the highly dense regions on the sky plane. This biases ${\rm\overline{{DM}}_{EoR}}$ relative to our initial results and also introduces more LoS dispersion, particularly at high redshifts where the FRB abundance drops rapidly. Taking realistic estimates of the FRB rate evolution with $z$, one requires $\lesssim 100$ FRBs to be distributed across the whole EoR window in order to discern the reionization histories at $3\sigma$ (see the left column of Figure 16) using the mean DM only. This is assuming we have removed contribution from lower redshifts ($z<6$) from the ${\rm DM}$ measurements which is an ‘optimistic’ case. Considering a ‘pessimistic’ case where the low redshift contribution is present, we find that the numbers could shoot as high as $200-600$ if we focus on the mid-reionization redshift bin (within uncertainty of $\Delta z=0.5$. Using the marginalized DM to discern between the reionization histories at $3\sigma$ might require roughly $\sim 1000$ FRBs in total during EoR (see right column of Figure 16).
@@
The numbers presented above corresponds to a particular choice of telescope sensitivity and FRB population models and are expected to vary if we change them, however we expect them to stay within the order of magnitude. We plan to include these effects gradually in our future work along the line. We have successfully demonstrated the potential of the derivative of the structure function $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ and of $\sigma_{\rm DM}$ as probes of the ionization bubble sizes along with the reionization history. The $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$, being a derivative, suffers from large variance, and we need a large number of FRBs ($N_{\rm FRB}^{\rm tot}\lesssim 100~{}000$) within range $6\leq z\leq 15$ to suppress the variance significantly down. The computation of $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ and their marginalization here considers a uniform distribution of the FRBs on the regular comoving grid. This is only done for convenience, and the derivatives of the structure function are well defined also when the sample is very uneven with $z$ and/or $\delta\theta$. In reality, FRBs should be associated with the galaxies which are generally clustered around the high-density peaks that gets ionized first. Therefore, it is highly probable to find more FRBs within the ionized regions, and that might help to compute the structure function in the vicinity of the ionized bubbles. We also plan to investigate more deeply into this estimator which will be very useful in probing the ionized regions around the sources.
@@ Acknowledgements
AKS, RG and SZ acknowledge support by the Israel Science Foundation (grant no. 255/18). AKS is also supported by National Science Foundation (grant no. 2206602). RG also acknowledges support from SERB, DST Ramanujan Fellowship no. RJF/2022/000141. PB is supported by a grant (no. 2020747) from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel, by a grant (no. 1649/23) from the Israel Science Foundation and by a grant (no. 80NSSC 24K0770) from the NASA astrophysics theory program. PK’s work is funded in part by an NSF grant AST-2009619 and a NASA grant 80NSSC24K0770.

### v2

@@ 5 Discussion
The dispersion introduced in the FRB pulse, while it travels through the ionized medium, can be used in probing the electron distribution along the LoS during EoR. We demonstrated the use of the sky-averaged $\rm DM$} and its derivatives to discern between the different reionization histories. Beyond this, we primarily aim to make use of the sky-averaged and the angular dispersion in the $\rm DM$} estimates to extract information on the ionization bubbles during the EoR. Using a toy model (Figure 1) of the binary ionization field (within the range $6\leq z\leq 15$), we see that the ${\rm\overline{{DM}}_{EoR}}(z)$ first increases (starting from the end of EoR) roughly up to the midpoint of the reionization, $z_{\rm mid}$, and then tend to saturate as the ionized regions decreases toward the initial stages of the reionization (Figure 2). The derivative $d{\rm\overline{{DM}}_{EoR}}/dz$ directly traces how fast the reionization progresses (Figure 3). The all-sky variance $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ is larger for the reionization scenario where the bubble sizes are larger and vice versa (Figure 6). We also compute $\partial^{2}\Xi/\partial z~\partial\delta\theta$ where $\Xi$ is the angular dispersion (eq. 4) at any redshift. We demonstrated that the contours are elongated or squeezed along $z$ and $\delta\theta$ depending respectively on the varying reionization history (Figure 4) and bubble sizes (Figure 7). The impacts are clearly prominent for the marginalized structure function derivatives (Figures 5 and 8).
@@
We also analyzed a more realistic reionization LC (Figure 10) generated using the grizzly 1D radiative transfer code for three different reionization histories ending at the same $z$. The behavior of $\rm\overline{{DM}}_{EoR}$} is the same as in the toy model. We find that the different LoS variance plays a significant role for realistic IGM (Figure 11) and biases the average ${\rm\overline{{DM}}_{EoR}}(z)$ by more than $1\sigma$ as compared to the case when the $\rm DM$} is computed using the averaged ionization fraction $\bar{x}_{\rm{H}\,\textsc{ii}}$. The slope of the $d{\rm\overline{{DM}}_{EoR}}/dz$ is directly sensitive to the reionization history (Figure 12). $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ is also dependent on the reionization history as well as indirectly on the IGM morphology (bubble sizes) in an intermingled way. Figure 14 clearly shows that $\partial^{2}\Xi/\partial z~\partial\delta\theta$ is sensitive to the reionization window as well as the typical bubble sizes. The contours are squeezed along the $z$-axis and elongated along the $\delta\theta$-axis for faster reionization history, which has relatively larger ionized bubbles and vice versa. Our initial analyses are rather optimistic, where we have considered the FRBs to be located uniformly at every grid point in our reionization LC with their redshifts known within an uncertainty of $\Delta z=0.5$. We also assumed that the contribution from the low redshift ($z<6$) has been perfectly removed from the DM measurements.
@@
We next consider an observationally more realistic situation where the FRB abundances vary with redshift (Figure 15) and they are more clustered at the highly dense regions on the sky plane. This biases $\rm\overline{{DM}}_{EoR}$} relative to our initial results and also introduces more LoS dispersion, particularly at high redshifts where the FRB abundance drops rapidly. Taking realistic estimates of the FRB rate evolution with $z$, one requires $\lesssim 100$ FRBs to be distributed across the whole EoR window in order to discern the reionization histories at $3\sigma$ (see the left column of Figure 16) using the mean DM only. This is assuming we have removed contribution from lower redshifts ($z<6$) from the $\rm DM$} measurements, which is an ‘optimistic’ case. Considering a ‘pessimistic’ case where the low redshift contribution is present, we find that the numbers could shoot as high as 200600200-600 if we focus on the mid-reionization redshift bin (within uncertainty of $\Delta z=0.5$). Using the marginalized DM to discern between the reionization histories at $3\sigma$ might require roughly $\sim 1000$ FRBs in total during EoR (see right column of Figure 16).
@@
The numbers presented above correspond to a particular choice of telescope sensitivity and FRB population models and are expected to vary if we change them; however, we expect them to stay within the order of magnitude. We plan to include these effects gradually in our future work along the line. We have successfully demonstrated the potential of the derivative of the structure function $\partial^{2}\Xi/\partial z~\partial\delta\theta$ and of $\sigma_{\rm DM}$ as probes of the ionization bubble sizes along with the reionization history. The $\partial^{2}\Xi/\partial z~\partial\delta\theta$, being a derivative, suffers from large variance, and we need a large number of FRBs ($N_{\rm FRB}^{\rm tot}\lesssim 100~000$) within the range $6\leq z\leq 15$ to suppress the variance significantly. The computation of $\partial^{2}\Xi/\partial z~\partial\delta\theta$ and their marginalization here considers a uniform distribution of the FRBs on the regular comoving grid. This is only done for convenience, and the derivatives of the structure function are well defined also when the sample is very uneven with $z$ and/or $\delta\theta$. In reality, FRBs should be associated with the galaxies that are generally clustered around the high-density peaks that get ionized first. Therefore, it is highly probable to find more FRBs within the ionized regions, and that might help to compute the structure function in the vicinity of the ionized bubbles. We also plan to investigate more deeply into this estimator, which will be very useful in probing the ionized regions around the sources.
@@ Acknowledgements
A.K.S., R.G. and S.Z. acknowledge support by the Israel Science Foundation (grant No. 255/18). A.K.S. is also supported by the National Science Foundation (grant No. 2206602). R.G. also acknowledges support from SERB, DST Ramanujan Fellowship No. RJF/2022/000141. P.B. is supported by a grant (No. 2020747) from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel, by a grant (No. 1649/23) from the Israel Science Foundation, and by a grant (No. 80NSSC 24K0770) from the NASA astrophysics theory program. P.K.’s work is funded in part by an NSF grant AST-2009619 and a NASA grant 80NSSC 24K0770.